OpenStudy (anonymous):

The slope of a line is -1/5, and the y-intercept is 5. What is the equation of the line written in general form? x - 5y - 25 = 0 x - 5y - 5 = 0 x + 5y - 25 = 0 ima be perty proud of myself if this is rite @texaschic101 x+5y-25=0 which is C

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@satellite73

OpenStudy (texaschic101):

y = -1/5x + 5 Multiply the equation by 5 5y = -x + 25 add x to both sides x + 5y = 25 subtract 25 from both sides x + 5y - 25 = 0 you got it...good job :)

OpenStudy (anonymous):

yes ive learned from the best.....0_0

OpenStudy (texaschic101):

lol....no, your just smart

OpenStudy (anonymous):

lol im not smart why do you think im here